### v2

# MD 20742, USA

Abstract.
We study heavy tailed random variables obtained by sampling an observable with a power singularity along an orbit of an expanding map on the circle. In the regime where the observables have infinite variance, we show that the set of limit points of the ergodic sum process, rescaled by $N^{1/s}(\operatorname{ln}N)^{\alpha}$, where $s$ is the index of limiting stable law, equals to the set of piecewise constant increasing functions with finitely many jumps. The proof relies on the multiple Borel-Cantelli Lemma proven in Dolgopyat, Fayad and Liu [J. Mod. Dyn. 18 (2022) 209–289].
1. Introduction
Let $X_{k}$ be a sequence of independent and identically distributed random variables with mean 0 and variance 1. Let $S_{n}=\sum_{k=1}^{n}X_{k}.$ Consider a function $W_{N}:[0,1]\to\mathbb{R}$ defined by $W_{N}\left(\frac{n}{N}\right)=\frac{S_{n}}{\sqrt{N}}$ at rational points $\frac{n}{N}$ and by linear interpolation in between. The functional central limit theorem states that the distribution of $W_{N}(\cdot)$ converges to the standard Wiener measure as $N\to\infty$. Since Wiener measure has full support, it follows (see e.g. [41]) that with probability $1$ for any continuous function $h$ on $[0,1]$, there is a sequence $N_{j}$ such that $W_{N_{j}}$ converges uniformly to $h.$
It is natural to ask whether one can obtain a smaller set of limits by considering a larger normalization (so that the renormalized function is close to 00 most of the time). Strassen [42] answered this question by showing that if one takes
${\widetilde{W}}_{N}\left(\frac{n}{N}\right)=\frac{S_{n}}{\sqrt{2N\operatorname{ln}\operatorname{ln}N}},$
then with probability $1$, $h$ is a limit point of $\{{\widetilde{W}}_{N}\}$ if and only if $h$ is an absolutely continuous function such that $h(0)=0$ and $\int_{0}^{1}(h^{\prime}(x))^{2}{\rm{d}}x\leq 1.$ This result is referred to as the functional form of law of the iterated logarithm (FLIL) in the literature [16, 39]. FLIL is a generalization of Khintchine-Kolmogoroff’s original form [28, 29] of the law of the iterated logarithm (LIL) and Hartman–Wintner’s LIL [17] for random walks. The method used in [42] for proving FLIL is known as Strassen’s almost sure invariance principle (ASIP). ASIP was further investigated for martingales (see e.g. [16, 20, 43]), which states that a martingale satisfying certain moment conditions is with probability $1$ close to a Brownian motion. Consequently, FLIL holding for Brownian motion is still valid for martingales.
ASIP can be extended to dynamical systems [39] by approximating weakly dependent random variables by martingales. This method has been applied to show ASIP for various dynamical systems (see e.g. [7, 8, 13, 15, 18, 19, 21, 30, 31, 33, 36, 38, 40]). Specifically, these systems include uniformly hyperbolic systems [10], Lorenz attractors [22], non-uniformly hyperbolic systems [32, 34, 35], partially hyperbolic systems [11], dispersing billiards [3, 5] and compact group extensions [14]. However, there are still many examples of dynamical systems where the limit distributions of ergodic sums are non-Gaussian. For these systems, the above approach is not applicable and ASIP does not hold. It is therefore of interest to develop a new method for finding a suitable normalizing sequence $L_{N}$ so that the function ${\tilde{S}}_{N}$ defined by ${\tilde{S}}_{N}(\frac{n}{N})=\frac{S_{n}}{L_{N}}$ has a non-trivial set of limit points.
In this paper, we address the above problem in a simple setting. Let $S_{n}(y)=\sum_{k=1}^{n}\phi(f^{k}y)$ where $f$ is an expanding map of the circle $\mathbb{S}$ and $\phi$ has a power singularity of the type $|y-x|^{-1/s}$ for fixed $x\in\mathbb{S}.$ Assuming that $0<s<2,$ $s\neq 1$ and $x$ is a Lebesgue typical point, we show that a reasonable choice of $L_{N}$ is $L_{N}=N^{1/s}(\operatorname{ln}N)^{\alpha}.$ If $\alpha$ satisfies ${[(r+1)s]}^{-1}<\alpha\leq{(rs)}^{-1},$ then the set of limit points consists of piecewise constant increasing functions starting at 00 with at most $r$ jumps.
Before formulating our result precisely, let us mention two related topics.
The first topic is Laws of Large Numbers for trimmed sums (see [1, 23, 24, 25, 27, 37]). Trimming is an important tool in our analysis since it allows us to discard all the summands in $S_{N}$ except for $r$ largest ones. However, this is only part of our analysis, since the contribution of the retained terms needs to be analyzed in more detail than in the papers mentioned above. We also note that unlike the above mentioned papers, our results apply not only to sums with infinite mean but also to the sums with infinite variance.
The second topic is the application of the classical Borel-Cantelli Lemma to estimate the rate of convergence in the Law of Large Numbers (see [4, 6, 9, 26] and references wherein). This topic is closer to the subject of our paper. However, the papers mentioned above deal with a single $N$ at a time, while we obtain the functional version of the deviation result. Fundamental to our approach is the multiple Borel-Cantelli Lemma proposed recently in [12], which provides a criterion to determine when several rare events occur at the same time scale.
In Section 2, we present the main theorem and sketch its proof using the multiple Borel-Cantelli Lemma. In Section 3, we focus on a key lemma which allows us to control the trimmed sums. The proof in Section 3 uses the reflection principle for the processes under consideration. For completeness, we include some standard proofs as appendices.
2. An analogue of LIL for Smooth expanding maps
2.1. The main result.
Let $f$ be a $C^{2}$ expanding map of the circle $\mathbb{S}.$ Here we say $f$ is expanding if there exists $\lambda>1$ such that $|f^{\prime}(x)|>\lambda$ for any $x\in\mathbb{S}.$ Then there is a unique $f$-invariant measure $\mu$ that is equivalent to Lebesgue measure [44]. Fix $x,$ and let $\phi$ be an unbounded function of the form
$\phi(y)=a|y-x|^{-1/s}+\psi(y),$
where $\psi$ is a $C^{1}$ function and $s>0.$ Let
$X_{n}(y)=\phi(f^{n}y).$
Denote $\|\cdot\|_{\infty}$ the uniform norm and $L=\|\psi\|_{\infty}.$ Then $\mu(X_{n}>t)=\mu(\phi(y)>t)$ and
(2.1) $\mu\left(|y-x|<\left(\frac{a}{t+L}\right)^{s}\right)\leq\mu(\phi(y)>t)\leq\mu\left(|y-x|\leq\left(\frac{a}{t-L}\right)^{s}\right)$
for $t>L.$ It follows that
(2.2) $\mu(X_{n}>t)\sim a_{s}t^{-s}$
for sufficiently large $t$ and $a_{s}=\rho_{0}a^{s}$, where $\rho_{0}$ is the density of $\mu$ at $x.$ For a positive $\alpha$, set
$W_{N}(t)=\frac{\left(\sum_{n\leq Nt}X_{n}\right)-a_{N}}{N^{1/s}(\operatorname{ln}N)^{\alpha}},$
where $a_{N}=0$ if $s\leq 1$ and $a_{N}=Nt\int\phi(y){\rm{d}}\mu(y)$ if $s>1.$
We regard $W_{N}(\cdot)$ as an element of the space ${\mathbb{D}}[0,1]$ of functions from $[0,1]$ to [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] that are right continuous and have left limits, where we endow [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] with the distance $d(z_{1},z_{2})=\left|\tan^{-1}(z_{1})-\tan^{-1}(z_{2})\right|.$ There are several possible topologies on ${\mathbb{D}}[0,1]$, the most relevant for us is Skorokhod’s $J_{1}$ topology. Let $\Lambda$ be the space of order preserving homeomorphisms on $[0,1].$ Then $J_{1}$ topology is defined using the distance
$d_{J_{1}}(h_{1},h_{2})=\inf_{\lambda\in\Lambda}\max\left\{\|\lambda-Id\|_{\infty},\,\max_{t\in[0,1]}d(h_{1}(t),h_{2}(\lambda(t)))\right\}.$
Two functions are $J_{1}$ close, if their graphs are close with respect to the Hausdorff distance.
In our analysis we shall use several facts about $J_{1}$ topology. First, if two functions are close with respect to the supremum norm, then they are $J_{1}$ close. Second, if one takes ${\mathbb{H}}_{r}$ to be the space of step functions with (at most) $r$ jumps, then $W\in{\mathbb{H}}_{r}$ can be written in the form
$W(t)=\sum_{m=1}^{\bar{r}}c_{m}{\mathbf{1}}_{t\leq t_{m}}(t),$
where $\bar{r}\leq r$ and $\{c_{m}\}_{m=1}^{r}$ and $\{t_{m}\}_{m=1}^{r}$ are $\bar{r}$–tuples such that $c_{m}$’s are positive and $t_{m}$’s are increasing. Two elements $W_{1}$ and $W_{2}$ of ${\mathbb{H}}_{r}$ with exactly $r$ jumps are $J_{1}$ close provided that $\{c_{m}^{1}\}_{m=1}^{r}$ is close to $\{c_{m}^{2}\}_{m=1}^{r}$ and $\{t_{m}^{1}\}_{m=1}^{r}$ is close to $\{t_{m}^{2}\}_{m=1}^{r}$. Third, $J_{1}$ topology allows us to ignore small jumps.
We also need the following definition. We call $x$ Diophantine if there exists $\epsilon=\epsilon(x)>0$ and $\rho_{0}=\rho_{0}(x)>0$ such that for all positive $\rho\leq\rho_{0}$ and all positive integer $k\leq\epsilon|\operatorname{ln}\rho|,$ $f^{-k}B(x,\rho)\cap B(x,\rho)=\emptyset.$ Due to Lemma 5 in [2], $\mu-$almost every $x$ is Diophantine.
Theorem 2.1.
Suppose that either
(i) $10<s<$ 1 and $x$ is non-periodic or
(ii) $21\leq s<$ 2 and $x$ is Diophantine.
Let $\alpha$ be such that $\frac{1}{(r+1)s}<\alpha\leq\frac{1}{rs}$ for some $r\in{\mathbb{N}}.$ Then for almost every $y$ the following holds: the set of $J_{1}$ limit points of $\{W_{N}(t)\}_{N=1}^{\infty}$ is equal to ${\mathbb{H}}_{r}.$
2.2. Multiple Borel-Cantelli Lemma.
The proof of Theorem 2.1 requires control on the number of times $k\leq n$ such that the events $\left\{y:\,c_{j}-\varepsilon\leq X_{k}(y)\leq c_{j}+\varepsilon\right\}$ occur for some $c_{j}>0$ and $0<\varepsilon<c_{j}.$ If several such events occur for some $n$ we also need to control the occurrence time of these events. To this end, we shall use a generalization of the multiple Borel-Cantelli Lemma proven in [12], those statement is recalled in this subsection.
Consider a sequence $\{\rho_{n}\}_{n=1}^{\infty}$ and a family of events $\{\Omega_{\rho_{n}}\}_{n=1}^{\infty}.$ Suppose that there exist ${\hat{\varepsilon}}>0$ and a decomposition of $\Omega_{\rho_{n}}$ for each $\rho_{n}$ such that
$\Omega_{\rho_{n}}=\bigcup_{i=1}^{p}\Omega_{\rho_{n},i}\quad\text{and}\quad\mu(\Omega_{\rho_{n},i})\geq{\hat{\varepsilon}}\mu(\Omega_{\rho_{n}}).$
We denote by
(2.3) $\Omega^{k}_{\rho_{n}}=f^{-k}(\Omega_{\rho_{n}}),\,\,\,\Omega^{k}_{\rho_{n},i}=f^{-k}(\Omega_{\rho_{n},i}),\,\,\,\sigma_{\rho_{n}}=\mu(\Omega_{\rho_{n}})\,\,\,{\text{and}}\,\,\,\sigma_{\rho_{n},i}=\mu(\Omega_{\rho_{n},i}).$
Fix $r\in\mathbb{N^{*}},$ for each $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$ we will characterize whether with probability 1, for infinitely many $n,$ the events $\Omega_{\rho_{n},i_{j}}^{k_{j}}$ happen for all $1\leq j\leq r.$
For an arbitrary $r$-tuple $0<k_{1}<\dots<k_{r}\leq n,$ we take the separation indices
$\operatorname{Sep}_{n}(k_{1},\dots,k_{r})=\text{Card}\left\{j\in\{0,\dots r-1\}:k_{j+1}-k_{j}\geq s(n)\right\},\quad k_{0}:=0;$
$\widehat{{\rm Sep}}_{n}(k_{1},\dots,k_{r})=\text{Card}\left\{j\in\{0,\dots r-1\}:k_{j+1}-k_{j}\geq\hat{s}(n)\right\},\quad k_{0}:=0$
for functions $s:{\mathbb{N}}\rightarrow{\mathbb{N}},$ $s(n)\leq(\operatorname{ln}n)^{2}$ and $\hat{s}:{\mathbb{N}}\rightarrow{\mathbb{N}},$ $\varepsilon n\leq\hat{s}(n)<n(1-q)/{(2r)}$ for some $0<q<1,$ $0<\varepsilon<(1-q)/{(2r)},$ respectively.
Suppose the function $\sigma.$ and $\sigma_{\cdot,i}$ defined in (2.3) are increasing and there is a sequence ${\varepsilon}_{n}\to 0$ satisfying some of the mixing conditions below.
  • $(M1)_{r}$
    If $0<k_{1}<k_{2}<\dots<k_{r}\leq n$ are such that $\operatorname{Sep}_{n}(k_{1},\dots,k_{r})=r$ and for each $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$
    $\left(\prod_{j=1}^{r}\sigma_{\rho_{n},i_{j}}\right)(1-{\varepsilon}_{n})\leq\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{n},i_{j}}\right)\leq\left(\prod_{j=1}^{r}\sigma_{\rho_{n},i_{j}}\right)(1+{\varepsilon}_{n});$
  • $(M2)_{r}$
    There exists $K>0$ such that if $0<k_{1}<k_{2}<\cdots<k_{r}\leq n$ are such that $\operatorname{Sep}_{n}(k_{1},\dots,k_{r})=m<r$, then
    $\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{n}}\right)\leq{\frac{K\sigma_{{\rho}_{n}}^{m}}{(\operatorname{ln}n)^{100r}}};$
  • $(M3)_{r}$
    If $0<k_{1}<\dots<k_{r}<l_{1}<\dots<l_{r}$ are such that $2^{i}<k_{\alpha}\leq 2^{i+1},2^{j}<l_{\beta}\leq 2^{j+1}$ for $1\leq\alpha,\beta\leq r$, $j-i\geq b$ for some constant $b\geq 2,$
    $\widehat{{\rm Sep}}_{2^{i+1}}(k_{1},\ldots,k_{r})=r,\quad\widehat{{\rm Sep}}_{2^{j+1}}(l_{1},\ldots,l_{r})=r,\quad l_{1}-k_{r}\geq\hat{s}(2^{j+1})$
    and for each $(i_{1},i_{2},\dots,i_{r}),\,(j_{1},j_{2},\dots,j_{r})\in\{1,\dots,p\}^{r},$
    $\mu\left(\left[\bigcap_{\alpha=1}^{r}\Omega^{k_{\alpha}}_{{\rho}_{2^{i}},i_{\alpha}}\right]\bigcap\left[\bigcap_{\beta=1}^{r}\Omega^{l_{\beta}}_{{\rho}_{2^{j}},j_{\beta}}\right]\right)\leq\left(\prod_{\alpha=1}^{r}\sigma_{\rho_{2^{i}},i_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\sigma_{\rho_{2^{j}},j_{\beta}}\right)(1+{\varepsilon}_{i}).$
We consider the events
$A_{*}^{n}=\{\exists\,0<k_{1}<k_{2}<\dots<k_{r}\leq n,\,\,\operatorname{Sep}(k_{1},\dots k_{r})=r,\,\,\,\Omega^{k_{j}}_{{\rho}_{n}}\text{ happens}\}$
and
$A_{I_{1},\dots,I_{r},i_{1},\dots i_{r}}^{n}=\{\exists\,k_{j},\,\,k_{j}/n\in I_{j},\,\Omega^{k_{j}}_{{\rho}_{n},i_{j}}\text{ happens}\}.$
We also let
$(M4)_{*}\quad\quad\lim_{n\to\infty}\frac{\mu(A_{*}^{n})}{n^{r}\sigma_{\rho_{n}}^{r}}=1.$
$(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}\quad\quad\lim_{n\to\infty}\frac{\mu(A_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}^{n})}{n^{r}\Pi_{j=1}^{r}\left(\sigma_{\rho_{n},i_{j}}|I_{j}|\right)}=1.$
Lemma 2.2.
Given a family of events $\{\Omega_{\rho_{n}}\}_{\rho_{n}>0}$ and their decomposition $\Omega_{\rho_{n}}=\bigcup_{i=1}^{p}\Omega_{\rho_{n},i}$, define
${\mathbf{S}}_{r}=\sum_{j=1}^{\infty}\left(2^{j}{\sigma_{{\rho}_{2^{j}}}}\right)^{r}.$
We assume that $\rho_{n_{1}}\leq\rho_{n_{2}}$ for $n_{1}\geq n_{2}.$
Suppose ${\mathbf{S}}_{r}<\infty$ and $\Omega_{{\rho}_{n_{1}}}\subset\Omega_{{\rho}_{n_{2}}}$ for $n_{1}\geq n_{2}:$
(a) If $(M1)_{r}$ and $(M2)_{r}$ are satisfied, then with probability $1$, we have that for large $n,$ the events $\Omega_{\rho_{n}}^{k}$ happen for at most $(r-1)$ times.
(b) If $(M4)_{*}$ holds, then with probability $1$, $A^{n}_{*}$ does not occur for large $n.$
Suppose ${\mathbf{S}}_{r}=\infty.$ For any $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r}$ and any intervals $I_{1},I_{2}\dots I_{r}\subset[0,1]:$
If
(c) $(M1)_{k}$, $(M2)_{k}$ for $k=r+1,\dots,2r$ as well as $(M3)_{r}$ are satisfied and $\hat{s}(n)/n<\min_{0\leq i,\,j\leq r,\,\,i\neq j}d(I_{i},I_{j}),$ where $I_{0}=0;$ or
(d) $(M1)_{r},$ $(M3)_{r}$ are satisfied and $(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}$ holds for $\hat{s}(n)/n<\min_{0\leq i,\,j\leq r,\,\,i\neq j}d(I_{i},I_{j}),$ where $I_{0}=0,$ then with probability $1$, there are infinitely many $n$ such that for some $k_{1}(n),k_{2}(n)\dots k_{r}(n)$ with $k_{j}(n)/n\in I_{j},$ the events $\Omega^{k_{j}}_{\rho_{n},i_{j}}$ occur.
Proof.
Part (a) is Theorem 2.6 in [12]. We consider the case ${\mathbf{S}}_{r}<\infty$ and $(M4)_{*}$ holds. Suppose $\{\delta_{n}\}$ is a sequence such that $\rho_{n}=\delta_{2n}$ and
$A_{*,2}^{n}=\{\exists\,0<k_{1}<k_{2}<\dots<k_{r}\leq 2n,\,\,\operatorname{Sep}(k_{1},\dots k_{r})=r,\,\,\,\Omega^{k_{j}}_{{\rho}_{n}}\text{ happens}\}.$
We obtain that
$\lim_{n\to\infty}\frac{\mu(A_{*,2}^{n})}{(2n)^{r}\sigma_{\delta_{2n}}^{r}}=1$
using $(M4)_{*}.$ Since $\sigma_{\rho_{n}}=\mu(\Omega_{\rho_{n}})=\mu(\Omega_{\delta_{2n}})=\sigma_{\delta_{2n}},$
$\lim_{n\to\infty}\frac{\mu(A_{*,2}^{n})}{n^{r}\sigma_{\rho_{n}}^{r}}=2^{r}.$
By taking $n=2^{m},$ we have that $\mu(A_{*,2}^{2^{m}})=O\left(2^{rm}{\sigma_{{\rho}_{2^{m}}}^{r}}\right)$ and
$\sum_{m=1}^{\infty}\mu(A_{*,2}^{2^{m}})=O\left({\mathbf{S}}_{r}\right)<\infty.$
It implies that $A_{*,2}^{2^{m}}$ does not occurs for infinitely many $m$’s with probability 1 using the classical Borel-Cantelli Lemma. Then Part (b) follows from the fact that $A_{*}^{n}\subset A_{*,2}^{2^{m}}$ when $2^{m}\leq n\leq 2^{m+1}.$
Part (c) is Theorem 2.12 in [12]. Finally, part (d) holds because in the proof of part (c) in [12], $(M1)_{k}$ and $(M2)_{k}$ for $k=r+1,\dots,2r$ are only used to verify $(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}$. $\square$
We now turn to this paper’s problem to consider the following targets:
$\Omega_{\rho_{N}}=\{y:\phi(y)>\rho_{N}\}$
and
$\Omega_{\rho_{N},j}=\left\{y:\frac{\phi(y)}{\rho_{N}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}]\right\}$
for $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{t},$ $t\in\mathbb{R}$ and $c_{j}>{\varepsilon}$ for any $j.$ Take $s(N)=K\operatorname{ln}N$ with $K$ being sufficiently large and $\hat{s}(N)=2{\varepsilon}N.$
Proposition 2.3.
(a) For any $r\in\mathbb{N},$ $\Omega_{\rho_{N}}$ and $\Omega_{\rho_{N},j}$ satisfy $(M1)_{r}$ and $(M3)_{r}.$
(b) If $x$ is Diophantine then $(M2)_{r}$ holds.
(c) If $x$ is non-periodic then $(M4)_{*}$ and $(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}$ are satisfied.
The proof of Proposition 2.3 is similar to Proposition 3.9 in [12] with some modifications, which is included in Appendix B.
Next, we will recall the exponential mixing property for expanding maps, which is useful not only in the proof Proposition 2.3 but also in the next section. Denote by $BV$ the space of functions with bounded variation on $\mathbb{S}$ and
${\mathbb{E}}\psi=\int\psi{\rm{d}}\mu\,\,\text{for}\,\,\psi\in BV.$
By Proposition 3.8 of [44], there are constants $C,$ $\theta<1$ such that
(2.4) $\left|{\mathbb{E}}\left(\psi_{1}(\psi_{2}\circ f^{n})\right)-({\mathbb{E}}\psi_{1})(E\psi_{2})\right|\leq C\|\psi_{1}\|_{BV}\|\psi_{2}\|_{L^{1}}\theta^{n},\,n\in\mathbb{N}$
for any $\psi_{1},\,\psi_{2}\in BV,$ where $\|\psi\|_{BV}:=\|\psi\|_{L^{1}}+\text{Var}(\psi)$ and $\text{Var}(\psi)$ is the variation of $\psi.$
In the following, we use $C$ to denote a constant that may change from line to line but that will not depend on $N.$
Lemma 2.4.
[…]
  1. (i)
    If $\varphi_{j}\geq 0$ for $1\leq j\leq q,$ we have for any $k_{1}\leq k_{2}\leq\cdots\leq k_{q},$ $k_{i}\in\mathbb{N},$
    (2.5) $\left(\prod_{j=1}^{q-1}\left(\|\varphi_{j}\|_{L^{1}}-C\|\varphi_{j}\|_{BV}\theta^{k_{j+1}-k_{j}}\right)\right)\|\varphi_{q}\|_{L^{1}}\leq\left\|\prod_{j=1}^{q}\varphi_{j}\circ f^{k_{j}}\right\|_{L^{1}}$
    $\leq\left(\prod_{j=1}^{q-1}\left(\|\varphi_{j}\|_{L^{1}}+C\|\varphi_{j}\|_{BV}\theta^{k_{j+1}-k_{j}}\right)\right)\|\varphi_{q}\|_{L^{1}}$
  2. (ii)
    If for each $j,$ $k_{j+1}-k_{j}\geq k,$ then
    (2.6) $\left|{\mathbb{E}}\left(\prod_{j=1}^{q}(\varphi_{j}\circ f^{k_{j}})\right)-\prod_{j=1}^{q}{\mathbb{E}}\varphi_{j}\right|\leq C\theta^{k}\prod_{j=1}^{q}\|\varphi_{j}\|_{BV}.$
Lemma 2.4 could be proved by the arguments of Proposition 6.2 in [12] with $\|\cdot\|_{Lip}$ replaced by $\|\cdot\|_{BV}.$ However, in the present case a simpler argument is available and we provide it in Appendix A for the convenience of the reader.
2.3. Proof of Theorem 2.1
Fix a small $\delta>0$ and let
$S_{n}^{\prime}=\sum_{1\leq k\leq n}\left(X_{k}{\mathbf{1}}_{\{X_{k}\geq N^{1/s}(\operatorname{ln}N)^{\delta}\}}-b_{N}^{\prime}\right),\quad S_{n}^{\prime\prime}=\sum_{1\leq k\leq n}\left(X_{k}{\mathbf{1}}_{\{X_{k}<N^{1/s}(\operatorname{ln}N)^{\delta}\}}-b_{N}^{\prime\prime}\right),$
where $b_{N}^{\prime}=b_{N}^{\prime\prime}=0$ if $s<1$ and $b_{N}^{\prime},b_{N}^{\prime\prime}$ are chosen to make the expectations of the LHS’s zero if $s>1.$
Lemma 2.5.
If $\delta<\alpha$ and either $10<s<$ 1 or $21<s<$ 2 and $x$ is Diophanitne then
[…]
The proof of Lemma 2.5 is the main part of this paper and is contained in the next section. Given Lemma 2.5, we proceed to show Theorem 2.1.
Let
$W_{N}^{\prime}(t)=\frac{S_{Nt}^{\prime}}{N^{1/s}(\operatorname{ln}N)^{\alpha}}.$
By Lemma 2.5, the limit points of $W_{N}(\cdot)$ and $W_{N}^{\prime}(\cdot)$ are the same. Consider the targets $\Omega_{\rho_{N}}=\{y:\phi(y)>\rho_{N}\}$ for $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{\delta}$ and suppose that $x$ is non-periodic. By Proposition 2.3 and Lemma 2.2(b), for large $N$ there are at most $\left[(\delta s)^{-1}\right]$ intervals $I_{j}$ of length $2s(N)$ such that all numbers $n$ such that $X_{n}>N^{1/s}(\operatorname{ln}N)^{\delta}$ are contained in one of those intervals. Moreover taking ${\bar{\alpha}}$ such that $((r+1)s)^{-1}<{\bar{\alpha}}<\alpha,$ we see that among those intervals there are at most $r$ which contain $n^{\prime}$s with $X_{n}\geq N^{1/s}(\operatorname{ln}N)^{{\bar{\alpha}}}.$ We claim that the contribution of the remaining intervals is negligible for a.e. x. Indeed suppose there is an interval $I$ of length $2s(N)$ and ${\bar{n}}\in I$ such that
(2.7) $\max_{n\in I}X_{n}=X_{\bar{n}}\leq N^{1/s}(\operatorname{ln}N)^{\bar{\alpha}}\;\;\mathrm{and}\;\;\sum_{n\in I}X_{n}\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}.$
Let
$I^{\prime}=\{n\in I:|n-{\bar{n}}|\leq(\operatorname{ln}N)^{(\alpha-{\bar{\alpha}})/2}\},\quad I^{\prime\prime}=I\setminus I^{\prime},\quad Y_{I}=\max_{n\in I^{\prime\prime}}X_{n}.$
Then
$\sum_{n\in I}X_{n}\leq\sum_{n\in I^{\prime}}X_{n}+\sum_{n\in I^{\prime\prime}}X_{n}\leq N^{1/s}(\operatorname{ln}N)^{(\alpha+{\bar{\alpha}})/2}+2Y_{I}s(N).$
By the definition of $I$ we have that $Y_{I}\geq{\bar{\varepsilon}}N^{1/s}(\operatorname{ln}N)^{\alpha-1}$ for some ${\bar{\varepsilon}}={\bar{\varepsilon}}({\varepsilon})>0.$ Accordingly if the interval $I$ satisfying (2.7), then there are two moments ${\bar{n}},{\hat{n}}\in[1,N]$ such that
(2.8) $(\operatorname{ln}N)^{(\alpha-{\bar{\alpha}})/2}\leq|{\bar{n}}-{\hat{n}}|\leq 2s(N),\quad\min\left\{X_{\bar{n}},\,X_{\hat{n}}\right\}>{\bar{\varepsilon}}N^{1/s}(\operatorname{ln}N)^{\alpha-1}.$
Thus the lemma below shows that the contribution of the intervals satisfying (2.7) is negligible. Its proof is available in Appendix B.
Lemma 2.6.
For a.e. $y$, (2.8) happens only for finitely many $N.$
It then follows that the limit points of $W_{N}^{\prime}(\cdot)$ have at most $r$ positive jumps. That is, all limit points are in ${\mathbb{H}}_{r}.$
To see that all points in ${\mathbb{H}}_{r}$ are realized, fix $r$-tuples $(c_{1},\dots,c_{r})$ and $(t_{1},\dots t_{r})$ such that $c_{j}>0$ for all $j$ and fix $0<{\varepsilon}<c_{j}.$ Consider the events
$\Omega_{\rho_{N},j}=\left\{y:\frac{\phi(y)}{\rho_{N}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}]\right\}.$
Taking $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{\alpha}$ and applying Proposition 2.3 and Lemma 2.2(d), we see that with probability $1$ there is a subsequence $N_{k}$ and times $n_{1}(k),n_{2}(k),\dots,n_{r}(k)$ such that
$\frac{n_{j}(k)}{N_{k}}\in[t_{j}-{\varepsilon},t_{j}+{\varepsilon}]\quad\text{and}\quad\frac{X_{n_{j}(k)}(y)}{N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}].$
It follows that for each $W\in{\mathbb{H}}_{r}$, $W$ belongs to the limit set with probability $1$ by taking a countable set of ${\varepsilon}_{l}$ converging to $0$ Recall the fact that two elements $W_{1}$ and $W_{2}$ of ${\mathbb{H}}_{r}$ with exactly $r$ jumps are $J_{1}$ close if and only if $\{c_{m}^{1}\}_{m=1}^{r}$ is close to $\{c_{m}^{2}\}_{m=1}^{r}$ and $\{t_{m}^{1}\}_{m=1}^{r}$ is close to $\{t_{m}^{2}\}_{m=1}^{r}$. Taking a countable dense subset in ${\mathbb{H}}_{r}$ completes the proof.
3. Proof of Lemma 2.5.
3.1. Splitting the sum
The proof of Lemma 2.5 will be given in the next three subsections. Fix a large number $D$ satisfying
(3.1) $D(1-s)+\alpha>1$ $\quad\text{if}\quad s<1;$
(3.2) $D(2-s)+2\alpha>3$ $\quad\text{if}\quad 1\leq s<2.$
and split $S_{N}^{\prime\prime}=\bar{S}_{N}+\bar{\bar{S}}_{N}$ where $\bar{S}_{n}=\sum_{k=1}^{n}\bar{X}_{k}$ and $\bar{\bar{S}}_{n}=\sum_{k=1}^{n}\bar{\bar{X}}_{k}$ for
$\bar{X}_{k}=X_{k}{\mathbf{1}}_{\{X_{k}<N^{1/s}(\operatorname{ln}N)^{-D}\}}-\bar{b}_{N}\,\,\text{and}\,\,\bar{\bar{X}}_{k}=X_{k}{\mathbf{1}}_{\{N^{1/s}(\operatorname{ln}N)^{-D}\leq X_{k}<N^{1/s}(\operatorname{ln}N)^{\delta}\}}-\bar{\bar{b}}_{N},$
where, as before, $\bar{b}_{N}=\bar{\bar{b}}_{N}=0$ for $s\leq 1$ and they are chosen to make the expectations of LHS’s zero for $s>1.$
We will show that
[…]
and
[…]
In § 3.2 we consider the case $s<1$ where the argument is much simpler since the sums ${\bar{S}}_{n}$ and ${\bar{{\bar{S}}}}_{n}$ are almost monotone. The case $21\leq s<$ 2 is more involved and requires the use of the reflection principle (see equations (3.14) and (3.23)). In § 3.3 and § 3.4 we prove (3.3) and (3.4) respectively for $21\leq s<$ 2 .
3.2. The case of infinite expectation.
Here we prove (3.3) and (3.4) in case $s<1.$
We begin with (3.3). By (2.2) we have
$E\left|\bar{S}_{N}\right|\leq N\int_{0}^{N^{1/s}(\operatorname{ln}N)^{-D}}\frac{a_{s}{\rm{d}}t}{t^{s}}+LN\leq CN^{1/s}(\operatorname{ln}N)^{-D(1-s)}.$
Hence
$E\left|\frac{\bar{S}_{N}}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\right|\leq\frac{C}{(\operatorname{ln}N)^{D(1-s)+\alpha}}.$
Take $N_{k}=2^{k}$, then for each $\varepsilon>0$,
(3.5) $\sum_{k\geq 1}\mu\left(\left|\frac{\bar{S}_{N_{k}}}{N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}}\right|>\varepsilon\right)\leq\sum_{k\geq 1}\frac{C}{\varepsilon k^{D(1-s)+\alpha}}<\infty$
where the convergence follows from (3.1). Hence Borel-Cantelli Lemma gives
[…]
Next, consider $2^{k}<N\leq 2^{k+1}.$ Since $|\bar{S}_{N}|\leq|\bar{S}_{N_{k+1}}|+2LN_{k+1}$ and $s>1$, (3.6) gives
[…]
proving (3.3).
Next, we prove (3.4). Choose $D$ such that $D=q\delta$ where $q$ is a positive integer. For $j=0,-1,\cdots,-q,$ let
$X_{k,j}=X_{k}{\mathbf{1}}_{\{X_{k}\in(N^{1/s}(\operatorname{ln}N)^{j\delta},N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}]\}},\quad{\bar{{\bar{S}}}}_{n,j}=\sum_{k=1}^{n}X_{k,j},$
where ${\mathbf{1}}.$ is the indicator function. It suffices to show that for each $j$,
(3.8) $\max_{1\leq n\leq N}\frac{|{\bar{{\bar{S}}}}_{n,j}|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\to 0.$
Since $|{\bar{{\bar{S}}}}_{n,j}|\leq|{\bar{{\bar{S}}}}_{N,j}|+2LN$ it suffices to prove that
(3.9) $\frac{|{\bar{{\bar{S}}}}_{N,j}|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\to 0.$
Next, we claim that for each $m$ there is a constant $C_{m}$ such that
(3.10) $E\left({\bar{{\bar{S}}}}_{N,j}^{m}\right)\leq C_{m}N^{m/s}(\operatorname{ln}N)^{\delta m}.$
Let ${\hat{B}}_{N,j}=\left\{y:\phi(y)\in\left(N^{1/s}(\operatorname{ln}N)^{j\delta},\;N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}\right]\right\},$ $j=0,-1,\cdots,-q.$ Note that
$\frac{{\bar{{\bar{S}}}}_{N,j}}{N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}}\leq\sum_{n=1}^{N}{\mathbf{1}}_{{\hat{B}}_{N,j}}(f^{n}y).$
Thus
(3.11) $E\left(\left(\frac{\bar{\bar{S}}_{N,j}}{N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}}\right)^{m}\right)\leq\sum_{1\leq k_{1}\leq\cdots\leq k_{m}\leq N}\mu\left(\bigcap_{j=1}^{m}\left\{f^{k_{j}}y\in{\hat{B}}_{N,j}\right\}\right).$
The next result is proven in Appendix B.
Lemma 3.1.
There is a constant $C$ such that
$\sum_{1\leq k_{1}\leq\cdots\leq k_{m}\leq N}\mu\left(\bigcap_{j=1}^{m}\left\{f^{k_{j}}y\in{\hat{B}}_{N,j}\right\}\right)\leq C\left(N\mu\left({\hat{B}}_{N,j}\right)\right)^{m}.$
Combining the lemma with (2.2) we see that (3.11) is less than $C(\operatorname{ln}N)^{-j\delta sm}.$ Since $j\leq 0$ and $s<1,$ (3.10) follows.
By Markov inequality
$\mu\left(|{\bar{{\bar{S}}}}_{N,j}|\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{C}{{\varepsilon}^{m}(\operatorname{ln}N)^{(\alpha-\delta)m}}$
for even $m.$ Taking $m$ such that $m(\alpha-\delta)>1$ and arguing as in the proof of (3.7) we obtain
[…]
Since $j$ is arbitrary, we obtain (3.4) for $s<1.$
3.3. Controlling small values of $X$ for $21\leq s<$ 2
Here we prove (3.3) in the case $21\leq s<$ 2 .
Given $y\in\mathbb{S}$, let $C_{n}(y)$ be the continuity component of $\phi 1_{\phi\leq N^{1/s}(\operatorname{ln}N)^{-D}}\circ f^{k},$ $k\leq n$ containing $y$ and $f^{n}|_{C_{n}(y)}$ be a diffeomorphism. Let $I_{n}(y)=f^{n}C_{n}(y).$ We say that $C_{n}(y)$ is short if the length of $I_{n}(y)$ is less than $N^{-1}(\operatorname{ln}N)^{-100}$ and we say that it is long otherwise. Fix $\varepsilon>0,$ we say that $C_{n}$ is extremal if there exists $z\in C_{n}$ such that $|{\bar{S}}_{n}(z)|\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}$ and for each $k<n,$ $|{\bar{S}}_{k}(z)|<{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}$ on $C_{n}.$ Note that different extremal components are disjoint. In addition,
(3.13) $|{\bar{S}}_{n}(y)|\geq\frac{3}{4}{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}\quad\text{for all }y\text{ in }C_{n}.$
Indeed, fix a large $K$ and write
$\left|{\bar{S}}_{n}(y)-{\bar{S}}_{n}(z)\right|\leq\sum_{k=1}^{n-K\operatorname{ln}N}|{\bar{X}}_{k}(y)-{\bar{X}}_{k}(z)|+|{\bar{S}}_{K\operatorname{ln}N}(f^{n-K\operatorname{ln}N}y)|+|{\bar{S}}_{K\operatorname{ln}N}(f^{n-K\operatorname{ln}N}z)|,$
then the last two terms are smaller than $2KN^{1/s}(\operatorname{ln}N)^{1-D}$ due to the cutoff in the definition of ${\bar{S}}_{n}.$ To estimate the first term, we note that since $d(f^{n}y,f^{n}z)\leq 1$ and $f$ is expanding then for $k\leq n-K\operatorname{ln}N,$ $d(f^{k}y,f^{k}z)\leq\theta^{K\operatorname{ln}N}.$ Since $\phi{\mathbf{1}}_{\{\phi<N^{1/s}(\operatorname{ln}N)^{-D}\}}$ is Lipschitz on its continuity components with Lipschitz constant less than $CN^{1+1/s}(\operatorname{ln}N)^{-(1+s)D}$, it follows that if $K$ is sufficiently large then $|{\bar{X}}_{k}(y)-{\bar{X}}_{k}(z)|\leq N^{-100}$, proving (3.13).
Let ${\mathcal{E}}_{N}$ be the set of points which belong to some extremal component for $n\leq N.$
Lemma 3.2.
If $D$ is large enough then $\mu({\mathcal{E}}_{N})\leq{C}{(\operatorname{ln}N)^{-2}}.$
Lemma 3.2 implies (3.3). In fact, combining this lemma with Borel-Cantelli Lemma we conclude that with probability $1$, for large $k$, $|{\bar{S}}_{n}|<{\varepsilon}N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}$ for $n\leq N_{k},$ $N_{k}=2^{k}.$ Since ${\varepsilon}$ is arbitrary, (3.3) follows.
Proof of Lemma 3.2.
Note that $\phi{\mathbf{1}}_{\{\phi<N^{1/s}(\operatorname{ln}N)^{-D}\}}$ is only discontinuous at two points $z_{-,N}$ and $z_{+,N}$ where $\phi(z_{\pm,N})=N^{1/s}(\operatorname{ln}N)^{-D}.$ Therefore, if $y$ belongs to a short component, there exists $k\leq n$ such that
$d\left(f^{k}y,\{z_{-,N},z_{+,N}\}\right)\leq N^{-1}(\operatorname{ln}N)^{-100}.$
Since for a fixed $k,$ $\mu\left(\left\{y:\,d\left(f^{k}y,\{z_{-,N},z_{+,N}\}\right)\leq N^{-1}(\operatorname{ln}N)^{-100}\right\}\right)=O\left(N^{-1}(\operatorname{ln}N)^{-100}\right),$ we obtain that the measure of points which belong to short components is $O\left((\operatorname{ln}N)^{-100}\right).$ It remains to estimate the measure of points which belong to long extreme components. Let $C_{n}$ be one such component and $I_{n}=f^{n}C_{n}.$ We will prove two estimates:
(3.14) $\mu\left(y\in C_{n}:|{\bar{S}}_{N}(y)|\geq\frac{{\varepsilon}}{2}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\geq C\mu(C_{n})$
for some $C>0$ independent of $N,$ $C_{n}$ and
(3.15) $\mu\left(y\in\mathbb{S}:|{\bar{S}}_{N}(y)|\geq\frac{{\varepsilon}}{2}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{C}{(\operatorname{ln}N)^{2}}$
for some $C>0$ independent of $N.$
To prove (3.15) we use Chebyshev inequality. By (2.2) we have
(3.16) ${\rm Var}({\bar{X}}_{n})\leq CN^{2/s-1}(\operatorname{ln}N)^{{-D(2-s)}}.$
We now estimate
(3.17) $\left|\text{Cov}({\bar{X}}_{n},{\bar{X}}_{m})\right|\leq\begin{cases}O\left(N^{2/s-1}(\operatorname{ln}N)^{{-D(2-s)}}\right)&\text{if }|m-n|<K\operatorname{ln}N,\\ N^{-2}&\text{if }|m-n|\geq K\operatorname{ln}N,\end{cases}$
where the first bound follows from (3.16) and Cauchy-Schwartz inequality and the second bound follows from (2.4).
Summing over all $m,n$ we get
(3.18) ${\rm Var}({\bar{S}}_{N})\leq C\left(N^{2/s}(\operatorname{ln}N)^{{1-D(2-s)}}+1\right).$
Since $D(2-s)+2\alpha>3$ due to (3.2), (3.15) follows by Chebyshev inequality.
The proof of (3.14) follows similar ideas. Namely, in view of (3.13), it suffices to show that there exists $C<1$ such that
(3.19) $\mu\left(y\in C_{n}:\left|\sum_{k=n+1}^{N}{\bar{X}}_{k}(y)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq C\mu(C_{n}).$
Since $f^{n}$ has bounded distortion property, (3.19) follows from
(3.20) $\mu\left(z\in I_{n}:|{\bar{S}}_{N-n}(z)|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq C\mu(I_{n}).$
Split ${\bar{S}}_{N-n}={\bar{S}}_{K\operatorname{ln}N}+({\bar{S}}_{N-n}-{\bar{S}}_{K\operatorname{ln}N}).$ The first term is bounded by
(3.21) $O\left(N^{1/s}(\operatorname{ln}N)^{1-D}\right).$
Note that
${\mathbb{E}}\left(({\bar{S}}_{N-n}-{\bar{S}}_{K\operatorname{ln}N}){\mathbf{1}}_{I_{n}}\right)^{2}=\sum_{K\operatorname{ln}N<k_{1}\leq k_{2}\leq N-n}{\mathbb{E}}\left(\bar{X}_{k_{1}}\bar{X}_{k_{2}}{\mathbf{1}}_{I_{n}}\right).$
Recall that ${\bar{X}}_{k_{1}}=\phi_{N}\circ f^{k_{1}}$ for $\phi_{N}=\phi{\mathbf{1}}_{\{\phi<N^{1/s}(\operatorname{ln}N)^{-D}\}}.$ Thus denoting $p=K\operatorname{ln}N/2,$ we obtain
${\mathbb{E}}\left(\bar{X}_{k_{1}}\bar{X}_{k_{2}}{\mathbf{1}}_{I_{n}}\right)={\mathbb{E}}\left((\phi_{N}\circ f^{k_{1}-p})(\phi_{N}\circ f^{k_{2}-p})\right)\mu(I_{n})+O\left(\left\|(\phi_{N}\circ f^{k_{1}-p})(\phi_{N}\circ f^{k_{2}-p})\right\|_{L_{1}}\eta_{n,p}\right)$
by (2.4), where
$|\eta_{n,p}|\leq C\|I_{n}\|_{BV}\theta^{p}\leq C\theta^{p}.$
If $K$ and hence $p$ is large enough then $|\eta_{n,p}|\leq N^{-(2/s+100)}$ and so
$\sum_{K\operatorname{ln}N\leq k_{1}\leq k_{2}\leq N-n}\left\|(\phi_{N}\circ f^{k_{1}-p})(\phi_{N}\circ f^{k_{2}-p})\right\|_{L_{1}}\eta_{n,p}=O\left(N^{-98}\right).$
On the other hand, similarly to (3.18) we get
$\sum_{K\operatorname{ln}N\leq k_{1}\leq k_{2}\leq N-n}{\mathbb{E}}\left((\phi_{N}\circ f^{k_{1}-p})(\phi_{N}\circ f^{k_{2}-p})\right)\mu(I_{n})\leq C\left(N^{2/s}(\operatorname{ln}N)^{{1-D(2-s)}}+1\right)\mu(I_{n}).$
It follows that
$\int_{I_{n}}\left(\sum_{k=K\operatorname{ln}N}^{N-n}{\bar{X}}_{k}(z)\right)^{2}{\rm{d}}\mu(z)\leq C\mu(I_{n})\left[N^{2/s}(\operatorname{ln}N)^{1-D(2-s)}+1\right].$
Hence
(3.22) $\mu\left(z\in I_{n}:\left|\sum_{k=K\operatorname{ln}N}^{N-n}{\bar{X}}_{k}(z)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{16C\mu(I_{n})}{{\varepsilon}^{2}}(\operatorname{ln}N)^{1-D(2-s)-2\alpha}.$
Combining (3.21) and (3.22) we obtain (3.14). $\square$
3.4. Controlling intermediate values of $X$ for $21\leq s<$ 2
Here we prove (3.4) in case $21\leq s<$ 2 .
The proof of (3.4) for $s\geq 1$ will proceed similarly to §3.3. Namely we define extremal components similarly to §3.3 by replacing $\phi{\mathbf{1}}_{\{\phi<N^{1/s}(\operatorname{ln}N)^{-D}\}}$ with $\phi{\mathbf{1}}_{\{N^{1/s}(\operatorname{ln}N)^{-D}\leq\phi<N^{1/s}(\operatorname{ln}N)^{\delta}\}}$ and ${\bar{S}}_{n}$ with ${\bar{{\bar{S}}}}_{n}.$ We need to obtain the following analogues of (3.14) and (3.15) for long extremal components $C_{n}$:
(3.23) $\mu\left(y\in C_{n}:|{\bar{{\bar{S}}}}_{N}(y)|>\frac{{\varepsilon}}{2}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\geq C\mu(C_{n})$
for some $C$ independent of $N,$ $C_{n}$ and
(3.24) $\mu\left(y\in\mathbb{S}:|{\bar{{\bar{S}}}}_{N}(y)|\geq\frac{{\varepsilon}}{2}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{C}{(\operatorname{ln}N)^{2}}$
for some $C$ independent of $N.$
In addition, we claim that (3.13) still holds for almost every $y$ in $C_{n},$ even though the proof needs to be modified slightly. Namely it is not true that each summand in ${\bar{{\bar{S}}}}_{K\operatorname{ln}N}(y)$ is smaller than $N^{1/s}(\operatorname{ln}N)^{-D}.$ However, we claim that for almost every $y,$
[…]
for $N$ large enough. If (3.25) is not satisfied, then
(3.26) $\quad\quad\exists\,{\bar{n}},\,{\hat{n}}\leq K\operatorname{ln}N,\,\,|{\bar{n}}-{\hat{n}}|\geq(\operatorname{ln}N)^{(\alpha-\delta)/2},\,\min\left\{{\bar{{\bar{X}}}}_{\bar{n}},\,{\bar{{\bar{X}}}}_{\hat{n}}\right\}>{\bar{\varepsilon}}N^{1/s}(\operatorname{ln}N)^{-D}.$
The same argument of Lemma 2.6 shows that for almost every $y,$ (3.26) occurs for finitely many $N.$ Since the value of each term which is larger than $N^{1/s}(\operatorname{ln}N)^{-D}$ is smaller than $N^{1/s}(\operatorname{ln}N)^{\delta},$ we conclude that for almost every $y,$
(3.27) $\left|{\bar{{\bar{S}}}}_{K\operatorname{ln}N}(y)\right|\leq CN^{1/s}(\operatorname{ln}N)^{(\delta+\alpha)/2},$
which is sufficient to obtain (3.13).
Equation (3.24) follows from the Markov inequality and the following estimate.
Lemma 3.3.
Suppose $x$ is Diophantine. For each $m,$ we have
(3.28) $E\left({\bar{{\bar{S}}}}_{N}^{m}\right)\leq C_{m}N^{m/s}(\operatorname{ln}N)^{\delta m}.$
Proof.
Recall that we take
$\bar{\bar{X}}_{n}=X_{n}{\mathbf{1}}_{\{N^{1/s}(\operatorname{ln}N)^{-D}\leq X_{n}<N^{1/s}(\operatorname{ln}N)^{\delta}\}}-\bar{\bar{b}}_{N},\quad\bar{\bar{b}}_{N}=E\left(\phi{\mathbf{1}}_{\{N^{1/s}(\operatorname{ln}N)^{-D}\leq\phi<N^{1/s}(\operatorname{ln}N)^{\delta}\}}\right)$
for $21<s<$ 2 and
$\bar{\bar{X}}_{n}=X_{n}{\mathbf{1}}_{\{N^{1/s}(\operatorname{ln}N)^{-D}\leq X_{n}<N^{1/s}(\operatorname{ln}N)^{\delta}\}}$
for $s=1.$
Thus
(3.29) $E\left({\bar{{\bar{S}}}}_{N}^{m}\right)=\sum_{k=1}^{m}\sum_{1\leq n_{1}<\dots<n_{k}\leq N}\sum_{m_{1}+\dots+m_{k}=m}c_{m_{1},\dots,m_{k}}E(\bar{\bar{X}}_{n_{1}}^{m_{1}}\dots\bar{\bar{X}}_{n_{k}}^{m_{k}}),$
where $c_{m_{1},\dots,m_{r}}$ are combinatorial factors.
Note that by (2.2)
(3.30) ${\mathbb{E}}\left|\bar{\bar{X}}_{0}\right|^{l}\leq\left\{\begin{array}[]{ll}CN^{l/s-1}(\operatorname{ln}N)^{(s-l)D},&l<s,\\ CN^{l/s-1}(\operatorname{ln}N)^{(l-s)\delta},&l>s,\\ C\operatorname{ln}\operatorname{ln}N,&l=s,\\ \end{array}\right.$
and
$\left\|\bar{\bar{X}}^{l}_{0}\right\|_{BV}\leq CN^{l/s}(\operatorname{ln}N)^{l\delta}.$
(i) Suppose that $\operatorname{Sep}(n_{1},\dots,n_{k})=k.$ Then we use the multiple mixing bound of Lemma 2.4(ii). We take $K$ in the definition of separation condition so large that the error term in Lemma 2.4(ii) is $O(N^{-k}).$ There are two further subcases.
(a) $\min_{j}m_{j}>1.$ Using the second estimate in (3.30) we see that
$\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)\leq CN^{m/s-k}(\operatorname{ln}N)^{m\delta},$
while the error term is much smaller. Summing over all $n_{1},\dots,n_{k}$ we obtain this case gives the leading contribution of $CN^{m/s}(\operatorname{ln}N)^{m\delta}.$
(b) $\min_{j}m_{j}=1.$ If $1<s<2,$ then
$\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)=0,$
so each term is $O(N^{-k})$ and the total contribution of all terms is $O(1).$
If $s=1,$
$\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)$ $=$ $\left(\prod_{m_{j}\geq 2}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)\right)\left(\prod_{m_{j}=1}{\mathbb{E}}\bar{\bar{X}}_{n_{j}}\right)$
$\leq$ $C\left(\prod_{m_{j}\geq 2}N^{m_{j}-1}\left(\operatorname{ln}N\right)^{(m_{j}-1)\delta}\right)\left(\operatorname{ln}\operatorname{ln}N\right)^{m-k}\leq CN^{m-k}\left(\operatorname{ln}N\right)^{(m-k)\delta}\left(\operatorname{ln}\operatorname{ln}N\right)^{m}$
for $k<m$ and
$\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)=\prod_{j}{\mathbb{E}}\bar{\bar{X}}_{n_{j}}\leq C\left(\operatorname{ln}\operatorname{ln}N\right)^{m}$
for $k=m.$ It follows that the total contribution of all terms is bounded by $CN^{m}(\operatorname{ln}N)^{m\delta}$ when $\min_{j}m_{j}=1$ and $s=1.$
(ii) Suppose that $\operatorname{Sep}(n_{1},\dots,n_{k})=\bar{k}<k.$ Given that $x$ is Diophantine, we consider the case where
$\min_{j}\{n_{j+1}-n_{j}\}\geq\epsilon\operatorname{ln}N$
for some $\epsilon>0.$ Otherwise, the contribution to the sum is $0$ By (3.30),
${\mathbb{E}}\left|\bar{\bar{X}}_{n_{j}}\right|^{m_{j}}\leq CN^{m_{j}/s-1}(\operatorname{ln}N)^{m_{j}D}.$
In addition using Lemma 2.4(i), we see that $E(\bar{\bar{X}}_{n_{1}}^{m_{1}}\dots\bar{\bar{X}}_{n_{k}}^{m_{k}})$ is $O\left(N^{(m/s)-\bar{k}-\eta}(\operatorname{ln}N)^{Dm}\right),$ where $\eta=(k-\bar{k})\epsilon|\operatorname{ln}\theta|.$ Since the number of such terms in (3.29) is $O\left(N^{\bar{k}}(\operatorname{ln}N)^{k-\bar{k}}\right),$ we conclude that the contribution of these non-separated terms is $O\left(N^{m/s-\eta}(\operatorname{ln}N)^{(D+1)m}\right)$, which is much smaller than the RHS of (3.28). $\square$
The proof of (3.23) is similar to the proof of (3.14). Namely, as in Section 3.3, it suffices to have the following analogue of (3.20):
(3.31) $\mu\left(z\in I_{n}:|{\bar{{\bar{S}}}}_{N-n}(z)|>\frac{{\varepsilon}}{2}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq C\mu(I_{n}),$
for some $C>0.$ Split ${\bar{{\bar{S}}}}_{N-n}={\bar{{\bar{S}}}}_{K\operatorname{ln}N}+\left({\bar{{\bar{S}}}}_{N-n}-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}\right).$ The first term is estimated by (3.27). For the second term arguing in the same way as in the proof of (3.20) but replacing (3.18) by (3.28), we obtain
${\mathbb{E}}\left(\left({\bar{{\bar{S}}}}_{N-n}-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}\right)^{m}{\mathbf{1}}_{I_{n}}\right)\leq CN^{m/s}(\operatorname{ln}N)^{\delta m}\mu(I_{n}).$
Therefore,
(3.32) $\mu\left(z\in I_{n}:\left|{\bar{{\bar{S}}}}_{N-n}(z)-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}(z)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)$
$\leq\frac{16C\mu(I_{n})}{{\varepsilon}^{2}}(\operatorname{ln}N)^{-m(\alpha-\delta)}.$
Then we get (3.31) by combining (3.27) with (3.32).
Appendix A Exponential mixing of all orders for expanding maps
Proof of Lemma 2.4.
We prove part(i) by induction. When $q=2$ the statement reduces to (2.4). For $q>2$ we apply (2.4) with $\psi_{1}=\varphi_{1},$ $\psi_{2}=\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}-k_{2}})$ and obtain
$\left\|\prod_{j=1}^{q}\left(\varphi_{j}\circ f^{k_{j}}\right)\right\|_{L^{1}}=\|\varphi_{1}\|_{L^{1}}\|\psi_{2}\|_{L^{1}}+O\left(\theta^{k_{2}-k_{1}}\|\varphi_{1}\|_{BV}\|\psi_{2}\|_{L^{1}}\right),$
By induction
$\|\psi_{2}\|_{L^{1}}\leq\left(\prod_{j=2}^{q-1}\left(||\varphi_{j}||_{L^{1}}+\theta^{k_{j}-k_{j-1}}\|\varphi_{j}\|_{BV}\right)\right)\|\varphi_{q}\|_{L^{1}}.$
It follows that
$\left\|\prod_{j=1}^{q}\left(\varphi_{j}\circ f^{k_{j}}\right)\right\|_{L^{1}}\leq\left(\prod_{j=1}^{q-1}\left(\|\varphi_{j}\|_{L^{1}}+C\|\varphi_{j}\|_{BV}\theta^{k_{j+1}-k_{j}}\right)\right)\|\varphi_{q}\|_{L^{1}}.$
Proceeding in a similar way, we can prove that
$\left\|\prod_{j=1}^{q}\left(\varphi_{j}\circ f^{k_{j}}\right)\right\|_{L^{1}}\geq\left(\prod_{j=1}^{q-1}\left(\|\varphi_{j}\|_{L^{1}}-C\|\varphi_{j}\|_{BV}\theta^{k_{j+1}-k_{j}}\right)\right)\|\varphi_{q}\|_{L^{1}}.$
Now we turn to the proof of part (ii). When $q=2$ the statement also reduces to (2.4). Applying (2.4) with $\psi_{1}=\varphi_{1},$ $\psi_{2}=\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}-k_{2}}),$ we get
$\left|{\mathbb{E}}\left(\prod_{j=1}^{q}(\varphi_{j}\circ f^{k_{j}})\right)-\left({\mathbb{E}}\varphi_{1}\right){\mathbb{E}}\left(\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}})\right)\right|$
$\leq$ $C\theta^{k}\|\varphi_{1}\|_{BV}\left\|\left(\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}})\right)\right\|_{L_{1}}\leq C\theta^{k}\|\varphi_{1}\|_{BV}\left\|\left(\prod_{j=2}^{q}(|\varphi_{j}|\circ f^{k_{j}})\right)\right\|_{L_{1}}\leq C\theta^{k}\prod_{j=1}^{q}\|\varphi_{j}\|_{BV},$
where the last inequality is due to part (i) of the lemma, which we have already proved. Applying inductive estimate to ${\mathbb{E}}\left(\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}})\right)$ we obtain (2.6). $\square$
Appendix B Good targets for expanding maps
Proof of Proposition 2.3.
We prove for $\Omega_{\rho_{N},j},$ and the proof for $\Omega_{\rho_{N}}$ follows in the same approach. Denote $B(x,r_{1},r_{2})=\{y\in\mathbb{S}:r_{1}\leq|y-x|\leq r_{2}\},$ and $\tilde{A}^{+}_{\rho_{N},j},$ $\tilde{A}^{-}_{\rho_{N},j}$ to be the characteristic functions of
$B\left(x,\left(\frac{a}{\rho_{N}(c_{j}+\varepsilon)+L}\right)^{s},\left(\frac{a}{\rho_{N}(c_{j}-\varepsilon)-L}\right)^{s}\right)$
and
$B\left(x,\left(\frac{a}{\rho_{N}(c_{j}+\varepsilon)-L}\right)^{s},\left(\frac{a}{\rho_{N}(c_{j}-\varepsilon)+L}\right)^{s}\right),$
respectively. Then
$\tilde{A}^{-}_{\rho_{N},j}\leq{\mathbf{1}}_{\Omega_{\rho_{N},j}}\leq\tilde{A}^{+}_{\rho_{N},j},\quad E\tilde{A}^{+}_{\rho_{N},j}-E\tilde{A}^{-}_{\rho_{N},j}\leq C\rho_{N}^{-(s+1)}\text{ and }\|\tilde{A}^{*}_{\rho_{N},j}\|_{BV}\leq 2\mu(\Omega_{\rho_{N},j})+4,$
[…]
First we prove $(M1)_{r}.$ If $0<k_{1}<k_{2}<\dots<k_{r}\leq N$ are such that $k_{i+1}-k_{i}\geq K\operatorname{ln}N,$ where $K$ is a sufficiently large constant, and each $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$
$\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right)\leq{\mathbb{E}}\left(\prod_{j=1}^{r}\tilde{A}^{+}_{\rho_{N},i_{j}}\circ f^{k_{j}}\right)\leq$
$\left(\prod_{j=1}^{r-1}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{j}}+C\theta^{K\operatorname{ln}N}(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{j}}+4)\right)\right)({\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{r}})\leq\left(\prod_{j=1}^{r}\sigma_{\rho_{N},i_{j}}\right)(1+{\varepsilon}_{N})$
for ${\varepsilon}_{N}\rightarrow 0,$ which yields the RHS of $(M1)_{r}.$ The LHS is proved similarly replacing $\tilde{A}^{+}_{\rho_{N},i_{j}}$ by $\tilde{A}^{-}_{\rho_{N},i_{j}}$.
Second we show $(M2)_{r}$ assuming that $x$ is Diophanitine. Let $\operatorname{Sep}(k_{1},\dots k_{r})=m<r.$ Since $x$ is Diophantine, there exists $\epsilon>0$ such that $B(x,c\rho_{N}^{-s})\cap f^{-k}B(x,c\rho_{N}^{-s})=\emptyset$ for $k\leq\epsilon\operatorname{ln}N$ and $c=2a^{s}/c_{j}^{s}.$ Thus we assume that $k_{j+1}-k_{j}\geq\epsilon\operatorname{ln}N$ for all $j.$ Then
$\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right)$$\leq$ $C\left(\prod_{k_{j+1}-k_{j}>K\operatorname{ln}N}\sigma_{\rho_{N},i_{j}}\right)(\theta^{\epsilon\operatorname{ln}N})^{r-m}(1+{\varepsilon}_{N})\leq{\frac{C\sigma_{{\rho}_{N}}^{m}}{(\operatorname{ln}N)^{100r}}}.$
Third we prove $(M3)_{r}.$ If $k_{1}<\dots<k_{r}<l_{1}<\dots<l_{r}$ are such that $2^{i}<k_{\alpha}\leq 2^{i+1},2^{j}<l_{\beta}\leq 2^{j+1}$ for $1\leq\alpha,\beta\leq r$, then for each $(i_{1},i_{2},\dots,i_{r}),\,(j_{1},j_{2}\dots j_{r})\in\{1,\dots,p\}^{r},$
$\mu\left(\left[\bigcap_{\alpha=1}^{r}\Omega^{k_{\alpha}}_{{\rho}_{2^{i}},i_{\alpha}}\right]\bigcap\left[\bigcap_{\beta=1}^{r}\Omega^{l_{\beta}}_{{\rho}_{2^{j}},j_{\beta}}\right]\right)\leq{\mathbb{E}}\left(\left(\prod_{\alpha=1}^{r}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}\circ f^{k_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}\circ f^{k_{\beta}}\right)\right)$
$\leq\left(\prod_{\alpha=1}^{r}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}+C\theta^{Ki}(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}+4)\right)\right)\left(\prod_{\beta=1}^{r-1}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}+C\theta^{Kj}(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}+4)\right)\right)({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{r}})$
$\leq\left(\prod_{\alpha=1}^{r}\sigma_{\rho_{2^{i}},i_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\sigma_{\rho_{2^{j}},j_{\beta}}\right)(1+{\varepsilon}_{i}).$
Finally we prove $(M4)_{*}$ assuming that $x$ is non-periodic. The proof of $(M4)_{I_{1},\dots I_{r},i_{1},\dots,i_{r}}$ is similar and will be left to the reader. We denote by
$\tau_{*}^{N}=\sum_{\overset{0<k_{1}<k_{2}<\dots<k_{r}\leq N}{{\operatorname{Sep}_{n}}(k_{1},\dots k_{r})=r}}\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N}}\right).$
Using $(M1)_{r},$ which has been proved, we have $\lim_{N\rightarrow\infty}\tau_{*}^{N}/\left(N\sigma_{\rho_{N}}\right)^{r}=1.$ Therefore, it suffices to show that
$\lim_{N\rightarrow\infty}\frac{\mu(A_{*}^{N})}{\tau_{*}^{N}}=1.$
By Boole’s inequality $\mu(A_{*}^{N})\leq\tau_{*}^{N}.$ For the lower bound we note that
(B.1) $\mu\left(A_{*}^{N}\right)\geq\tau_{*,b}^{N}-\sum^{*}\mu(\Omega^{k_{1}}_{{\rho}_{N}}\cap\dots\cap\Omega^{k_{r}}_{{\rho}_{N}}\cap\Omega^{k_{r+1}}_{{\rho}_{N}}),$
where the sum in $\sum^{*}$ is taken over all tuples satisfying $\operatorname{Sep}(k_{1},\dots k_{r})=r.$ We now consider several cases depending on the size of $p=\min_{j=1,\dots,r}|k_{r+1}-k_{j}|$.
(i) Since $x$ is non-periodic for each fixed $p_{0},$ we have that $\Omega_{{\rho}_{N}}\cap f^{-p}\Omega_{{\rho}_{N}}=\emptyset$ for all $p\leq p_{0}$ provided that $n$ is large enough. Hence the contribution of terms $p\leq p_{0}$ is zero.
(ii) Next we suppose that $p_{0}<p\leq\varepsilon\operatorname{ln}N.$ Precisely we may assume that $k_{r+1}=k_{\bar{j}}+p$ for some ${\bar{j}}\leq r$ (the case where $k_{r+1}=k_{\bar{j}}-p$ is similar). Define $\bar{A}_{{\rho}_{N}}^{+}$ to be the characteristic function of $B\left(x,\left(\frac{a}{\rho_{N}-L}\right)^{s}\right).$ Then
$\bar{A}^{+}_{\rho_{N}}\geq{\mathbf{1}}_{\Omega_{\rho_{N}}},\,{\mathbb{E}}\bar{A}^{+}_{\rho_{N}}-\mu(\Omega_{\rho_{N}})\leq C\rho_{N}^{-(s+1)}\text{ and }\|\bar{A}^{+}_{\rho_{N}}\|_{BV}\leq 2\mu(\Omega_{\rho_{N}})+4.$
For $j=1,\dots,r,$ define
${\hat{A}}_{j}=\begin{cases}\bar{A}_{{\rho}_{N}}^{+}&\text{if }j\neq{\bar{j}},\\ \bar{A}_{{\rho}_{N}}^{+}(\bar{A}_{{\rho}_{N}}^{+}\circ f^{p})&\text{if }j={\bar{j}}.\end{cases}$
If we denote by $M=\max_{x\in\mathbb{S}}|f^{\prime}(x)|,$ then
$\mu\left(\prod_{j=1}^{r+1}\Omega_{{\rho}_{N}}^{k_{j}}\right)\leq{\mathbb{E}}\left(\prod_{j=1}^{r}{\hat{A}}_{j}\circ f^{k_{j}}\right)$
$\leq$ $\left(\prod_{j\neq{\bar{j}}}\left({\mathbb{E}}{\hat{A}}_{j}+C\|\bar{A}_{{\rho}_{N}}^{+}\|_{BV}\theta^{K\operatorname{ln}N}\right)\right)\left({\mathbb{E}}{\hat{A}}_{\bar{j}}+C\|\bar{A}_{{\rho}_{N}}^{+}\|_{BV}^{2}M^{p}\theta^{K\operatorname{ln}N}\right)\leq\prod_{j=1}^{r}\left({\mathbb{E}}{\hat{A}}_{j}+n^{-10}\right)$
if $s(n)=K\operatorname{ln}N$ and $K$ is large enough. Noting that
${\mathbb{E}}({\hat{A}}_{\bar{j}})={\mathbb{E}}\left(\bar{A}_{{\rho}_{N}}^{+}(\bar{A}_{{\rho}_{N}}^{+}\circ f^{p})\right)\leq C\sigma_{{\rho}_{N}}\theta^{p}$
by (2.4), we have that each term with $p$ as in the case (ii) is at most $C\sigma_{{\rho}_{N}}^{r}\theta^{p}.$ Summing over all $r$ tuples and over all $p\in(p_{0},\varepsilon\operatorname{ln}N]$ yields a total contribution of $CN^{r}\sigma_{\rho_{N}}^{r}\theta^{p_{0}}$ for some constant $C$ independent of $p_{0}.$
(iii) The terms where ${\varepsilon}\operatorname{ln}N<p\leq K\operatorname{ln}N$ are $O(\sigma_{\rho_{N}}^{r+\eta})$ for some $\eta=\eta({\varepsilon})>0.$ Thus the total contribution of those terms are $O(N^{r}(\operatorname{ln}N)\;\sigma_{\rho_{N}}^{r+\eta}),$ which is also negligible.
(iv) The terms where $p>K\operatorname{ln}N$ for sufficiently large $K$ is $O(\sigma_{\rho_{N}}^{r+1})$ similarly to the proof of $(M1)_{r}.$ Thus the total contribution of those terms are $O((N\sigma_{\rho_{N}})^{r+1})$ which is negligible compared to $\tau_{*}^{N}.$
Summing the estimates of cases (i)–(iii) above we see that the second term in the RHS of (B.1) is negligible comparing to the first, which proves $(M4)_{*}.$square
Proof of Lemma 2.6.
Let
$H_{N,{\bar{\varepsilon}}}=\left\{y:\exists\,{\bar{n}},\,{\hat{n}}\leq N,\,\,(\operatorname{ln}N)^{(\alpha-{\bar{\alpha}})/2}\leq|{\bar{n}}-{\hat{n}}|\leq 2s(N),\,\,\min\{X_{\bar{n}},\,X_{\hat{n}}\}>{\bar{\varepsilon}}N^{1/s}(\operatorname{ln}N)^{\alpha-1}\right\},$
and
$\tilde{H}_{k,{\bar{\varepsilon}}}=\left\{y:\exists\,{\bar{n}},\,{\hat{n}}\leq 2^{k+1},\,\,k^{(\alpha-{\bar{\alpha}})/2}\leq|{\bar{n}}-{\hat{n}}|\leq 2s(2^{k+1}),\,\,\min\{X_{\bar{n}},\,X_{\hat{n}}\}>{\bar{\varepsilon}}2^{k/s}k^{\alpha-1}\right\}.$
Since $H_{N,{\bar{\varepsilon}}}\subset\tilde{H}_{k,{\bar{\varepsilon}}}$ for $2^{k}\leq N\leq 2^{k+1},$ it suffices to show that $\tilde{H}_{k,{\bar{\varepsilon}}}$ happens only for finitely many $k.$ Using (2.4) and similar method in the proof of Proposition 2.3 we conclude that
$\mu(\tilde{H}_{k,{\bar{\varepsilon}}})\leq\frac{C2^{k+1}(k+1)}{2^{k}k^{s(\alpha-1)}}\theta^{k(\alpha-{\bar{\alpha}})/2}=2Ck^{s(1-\alpha)+1}2^{k(\alpha-{\bar{\alpha}})\operatorname{ln}\theta/2}.$
Hence $\sum_{k}\mu(\tilde{H}_{k,{\bar{\varepsilon}}})<\infty$ and the result follows from the classical Borel-Cantelli Lemma. $\square$
Proof of Lemma 3.1.
Let $p=\min_{1\leq j\leq m-1}\{k_{j+1}-k_{j}\}.$ Since $x$ is non-periodic, for each fixed $p_{0},$ we have that ${\hat{B}}_{N,j}\cap f^{-p}{\hat{B}}_{N,j}=\emptyset$ for all $p\leq p_{0}$ provided that $N$ is large enough because $x$ is non-periodic. Thus the contribution of the terms with tuples satisfying $p\leq p_{0}$ is zero for large $N.$
We suppose that $\operatorname{Sep}(k_{1},\dots k_{m})=l<m.$ If $p_{0}<p\leq\varepsilon\operatorname{ln}N,$ Arguing as the case $(M4)_{*}$(ii) in the proof of Proposition 2.3, we conclude that each term is smaller than $C\mu({\hat{B}}_{N,j})^{l}\theta^{p}$ for fixed $p.$ Since the total number of terms is $O\left(N^{l}p^{m-l}\right),$ we have that total contribution of the terms with $p={\hat{p}}$ is $O\left(N^{l}{\hat{p}}^{m-l}\mu({\hat{B}}_{N,j})^{l}\theta^{\hat{p}}\right).$ The sum over tuples with ${\hat{p}}\in[p_{0},{\varepsilon}\operatorname{ln}N]$ gives a total contribution of $CN^{l}{p_{0}}^{m-l}\mu({\hat{B}}_{N,j})^{l}\theta^{p_{0}}$ for fixed $l$ and $p_{0}$ large enough (independent of $N$).
On the other hand if $\varepsilon\operatorname{ln}N<p\leq K\operatorname{ln}N,$ then similarly to the proof of the case $(M4)_{*}$(iii) we obtain that
$\mu\left(\bigcap_{j=1}^{m}\left\{f^{n_{j}}y\in{\hat{B}}_{N,j}\right\}\right)\leq C\mu({\hat{B}}_{N,j})^{l+\eta}.$
Summing over all $r$ tuples and over all $p\in(\varepsilon\operatorname{ln}N,K\operatorname{ln}N]$ yields a total contribution that is not greater than $C\left(N\mu({\hat{B}}_{N,j})\right)^{l}\mu({\hat{B}}_{N,j})^{\eta}\left(\operatorname{ln}N\right)^{m-l}.$
Finally, if $\operatorname{Sep}(k_{1},\dots k_{m})=m,$ similarly to the case $(M1)_{r}$ in the proof of Proposition 2.3, we obtain that
$\mu\left(\bigcap_{j=1}^{m}\left\{f^{k_{j}}y\in{\hat{B}}_{N,j}\right\}\right)\leq C\mu({\hat{B}}_{N,j})^{m}.$
Summing over all tuples $(k_{1},\dots k_{m})$ and using that $(N\mu({\hat{B}}_{N,j}))^{l}\leq C(N\mu({\hat{B}}_{N,j}))^{m}$ for $l<m$, we obtain the result. $\square$